\begin{thebibliography}{99}
\vskip\baselineskip
\vskip\baselineskip

\bibitem{Cle} Clemens H.: Curves on generic hypersurfaces. Ann. Sci. Ec. Norm. Sup. 
\textbf{19}(1986), 629–636.
\bibitem{Dem1} Demailly J.P.: Towards the Green-Griffiths-Lang Conjecture,   Conference “Analysis and Geometry”, Tunis, March 2014, in honor of Mohammed Salah Baouendi, ed. by A. Baklouti, A. El Kacimi, S. Kallel, N. Mir, Springer Proc. Math. Stat., 127, Springer-Verlag, (2015), 141-159.
\bibitem{Dem2} Demailly J.P.: Recent results on the Kobayashi and Green-Griffiths-Lang conjecture, 
 hal-01683413v2, (2018).
 \bibitem{Ein1} Ein L.: Subvarieties of generic complete intersections. Invent. Math. \textbf{94}(1988), 163-169.
  \bibitem{Ein2} Ein L.: Subvarieties of generic complete intersections. Subvarieties of generic complete intersections, II. Math. Ann. \textbf{289}(1991), 465-471.
  \bibitem{GG} Green M. and Griffiths P.A.: Two applications of algebraic geometry to entire holomorphic mappings. The Chern Symposium (1979); Proc. Internal. Sympos. Berkeley, CA, (1979); Springer- Verlag, New York (1980), 41-74.
\bibitem{Gri} Griffiths P.A.: Holomorphic mappings into canonical algebraic varieties. Ann. of Math. 
\textbf{98}(1971), 439-458.
\bibitem{Ho} H\"ormander L.:  An introduction to complex analysis in several variables, North-Holland, (1990).
\bibitem{Hu1} Hu P.C.: Holomorphic mappings between spaces of different dimensions I,
         Math. Z.  \textbf{214}(1993), 567-577;  II, Math. Z.
         \textbf{215}(1994), 187-193.
\bibitem{HY} Hu P.C. and Yang C.C.: Value distribution theory related to number theory, Birkh\"auser Verlag, (2006).
\bibitem{Lang86} Lang S.: Hyperbolic and Diophantine analysis, Bull. Amer.
        Math. Soc. (2), \textbf{14}(1986), 159-205.
\bibitem{Lu} Lu  S.S.Y.: On hyperbolicity and the Green-Griffiths conjecture for surfaces. Geometric Complex Analysis, ed. by J. Noguchi et al., World Scientific Publishing Co. (1996), 401-408.
\bibitem{L-Y} Lu, S.S.Y. and Yau, S.T.: Holomorphic curves in surfaces of general type. Proc. Nat. Acad. Sci. USA,  \textbf{87}(1990), 80-82.
       \bibitem{Nadel} Nadel A. M.:  Hyperbolic surfaces in $\mathbb{P}^3$, Duke Math. J.  \textbf{58}(1989), 749-771.
\bibitem{No} Noguchi J. and Winkelmann J.: Nevanlinna theory in several complex variables and Diophantine approximation,
 A series of comprehensive studies in mathematics, Springer, (2014).
\bibitem{ru} Ru M.: Nevanlinna theory its relation to diophantine approximation, World Scientific Publishing, (2001).
    \bibitem{Siu87} Siu Y.T.: Defect relations for holomorphic maps
        between spaces of different dimensions, Duke Math. J. 
        (1) , \textbf{55}(1987), 213-251.
\bibitem{SY} Siu Y.T. and Yeung S.K.:  A generalized Block's theorem and the hyperbolicity of the complement of an ample divisor in an abelian variety, Math. Anal. \textbf{306}(1996), 743-758.
      \bibitem{S2002} Siu Y.T.: Some recent transcendental techniques in algebraic and complex geometry, ICM, Vol. I, (2002), 439-448.
     \bibitem{Stoll} Stoll W.: Value distribution on parabolic spaces,
        Lecture Notes in Math. Springer-Verlag, \textbf{600}(1977).
    
\end{thebibliography}

\documentclass[11pt]{amsart}

\usepackage{amsmath, amsfonts, amssymb,amsthm}
\usepackage{amstext}
\usepackage{mathrsfs}

\usepackage{float,epsf,subfigure}
\usepackage[all,cmtip]{xy}
\usepackage{hyperref}
\usepackage{algorithm}
\usepackage{algorithmic}
\usepackage{mathtools}
\usepackage{float}
\usepackage{bm}
\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{cor}[theorem]{Corollary}
\newtheorem{conj}[theorem]{Conjecture}

\newtheorem{def-thm}[theorem]{Definition-Theorem}
\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{defi}[theorem]{Definition}

\newtheorem{theoremnn}{Theorem}

\theoremstyle{definition}

{

\begin{document}
\title[Algebraic degeneracy of holomorphic curves]{Algebraic degeneracy of holomorphic curves}
\author[X.J. Dong and P.C. Hu] { Xianjing Dong and Peichu Hu }

\address{School of Mathematics \\ China University of Mining and Technology \\ Xuzhou \\ 221116 \\ Jiangsu \\ P. R. China}
\email{xjdong05@126.com}
\address{Department of Mathematics \\ Shandong University  \\  Shandong, Jinan, 250100, P.R. China}
\email{pchu@sdu.edu.cn}


\subjclass[2010]{30D35; 32H30.} \keywords{Green-Griffiths; holomorphic curve; value distribution}
\date{}
\maketitle \thispagestyle{empty} \setcounter{page}{1}

\begin{abstract}
\noindent  We consider  the algebraic degeneracy of holomorphic curves from a point of  view of meromorphic vector fields.   Employing the notion of Jocabian sections introduced by W. Stoll, 
we establish   a 
Second Main Theorem type inequality.  As  consequences,   several algebraic degeneracy theorems are obtained for holomorphic curves into a  complex projective variety.
\end{abstract}

\vskip\baselineskip

\setlength\arraycolsep{2pt}

\section{Introduction}

Recall that a smooth complex projective variety $X$ is said to be of \emph{general type} if   $K_X$  (canonical line bundle) is \emph{pseudo ample} (cf. \cite{Lang86}). 
The well-known  Green-Griffiths conjecture (cf. \cite{GG, Lang86, S2002})
stipulates  that 
\begin{conj}\label{conj1} Let  $X$ be  a  smooth complex projective  variety of general type. 
Then $X$  contains  no  algebraically non-degenerate holomorphic curves. 
\end{conj}
A stronger version than Conjecture \ref{conj1} is   Green-Griffiths-Lang conjecture \cite{Lang86} which asserts  that 
if  $X$ is of general type,  then there is a proper algebraic subvariety $Z$ of $X$ such that
every  nonconstant holomorphic curve $f:\mathbb C\rightarrow X$ is contained in $Z.$ 
For more details about GG and GGL 
conjectures,  refer to Clemens \cite{Cle},  Demaily \cite{Dem1,Dem2}, Ein \cite{Ein1,Ein2}, Griffiths \cite{Gri},   Lu-Yau \cite{L-Y},   Siu \cite{Siu87, S2002},  
Siu-Yeung \cite{SY},   and see also  \cite{HY, Lu, Nadel, No,  ru}, etc.
 
In this paper,  we   consider  the algebraic degeneracy of holomorphic curves into complex projective varieties in a viewpoint of   meromorphic vector fields.
 The  technical line  is to establish a Second Main Theorem type inequality of Nevanlinna theory.

To begin with, we  introduce the main results of the paper, some notations will be illustrated later.  Let $X$ be a smooth complex projective variety with a holomorphic line bundle $E$ over $X.$
 Consider  $f:\mathbb C^m\rightarrow X,$ a  holomorphic mapping into $X$ with $\dim_{\mathbb C}X=n>m.$ Then, $f$ induces a tangent mapping $f_* : T_{\mathbb C^m}\rightarrow T_{X}$
 between holomorphic tangent bundles. The first main result of the paper  is  the following  Second Main Theorem type inequality
 \begin{theoremnn}[=Theorem \ref{mainthm}]\label{abc}
Let $L$ be an ample line bundle over $X.$
 Set
 $$\mathfrak Z_j=f_*\Big(\frac{\partial}{\partial z_j}\Big), \ \
 j=1,\cdots,m.$$
Let  $\mathfrak X_1,$ $\cdots,\mathfrak X_{n-m}$ be linearly almost-independent meromorphic vector fields over $X,$ and let $\mathfrak X_1\wedge \cdots\wedge \mathfrak X_{n-m}$ be
of  pole order at most $E,$ i.e.,  there exists a nonzero  holomorphic section
$t: X\rightarrow E$ such that $t\otimes \mathfrak X_1\wedge\cdots \wedge \mathfrak X_{n-m}$ is holomorphic.
Assume  that 
  $\mathfrak Z_{1f}\wedge\cdots\wedge \mathfrak Z_{mf}\wedge \varphi\not\equiv 0$
with  $\varphi=(t\otimes \mathfrak X_1\wedge\cdots \wedge \mathfrak X_{n-m})_f.$
If $f(\mathbb C^m)\nsubseteq F^{-1}_\varphi(0)$, then
\begin{equation*}
T_f(r,K_X)+ N_{f, {\rm Ram}}(r)\leq_{\rm exc} T_f(r,E)+ O\Big(\log^+T_f(r,L)+\log^+\log r\Big).
\end{equation*}
\end{theoremnn}
Theorem \ref{abc} yields  the following  degeneracy result
\begin{theoremnn}[=Corollary \ref{Cor1}] 
 Set
 $$\mathfrak Z_j=f_*\Big(\frac{\partial}{\partial z_j}\Big), \ \
 j=1,\cdots,m.$$
Let  $\mathfrak X_1,$ $\cdots,\mathfrak X_{n-m}$ be linearly almost-independent meromorphic vector fields over $X,$ and let $\mathfrak X_1\wedge \cdots\wedge \mathfrak X_{n-m}$ be
of  pole order at most $E.$ If $E^*\otimes K_X$ is pseudo ample,  then every 
holomorphic mapping $f:\mathbb C^m\rightarrow X$  is algebraically degenerate or  satisfied with the equation
  $$\mathfrak Z_{1f}\wedge\cdots\wedge \mathfrak Z_{mf}\wedge (\mathfrak X_{1}\wedge\cdots\wedge \mathfrak X_{n-m})_f= 0.$$
 \end{theoremnn}
To more precisely, if $\dim_{\mathbb C} X=2,$  then we obtain 
 \begin{theoremnn}[=Theorem \ref{ddc}]\label{tm3}
Assume that  $\dim_{\mathbb C}X=2.$ Let $\mathfrak X$ be a nonzero meromorphic vector field of  pole order at most $E$  over $X.$  If $E^*\otimes K_X$ is pseudo ample, 
 then $X$  contains  no  algebraically non-degenerate holomorphic curves. 
\end{theoremnn}

Finally, we  generalize a theorem of  Siu \cite{S2002} (or see Nadel \cite{Nadel}). 
Let $\mathscr D$ be a meromorphic connection on $T_X$ with
pole order at most $E,$ i.e., $t\otimes\mathscr D$ is holomorphic for a nonzero section $t\in
H^0(X,E).$ We obtain 
 \begin{theoremnn}[=Theorem \ref{Siu-thm-Nst21}]\label{sxx}
Let $f:\mathbb{C}\rightarrow X$ be a  
holomorphic curve.
 If $E^{-n(n-1)/2}\otimes K_X$ is pseudo ample, 
then  the image of $f$ is  contained in the pole divisor $(t)$
of  $\mathscr D.$ 
\end{theoremnn}
 
\section{Jacobian sections}

We   will generalize the notion of  \emph{Jacobian sections} introduced by Stoll (cf. \cite{HY, Stoll}),
which  is useful to the proofs of our 
 main theorems.
In this section,  $M,X$ denote   complex manifolds of complex dimension $m, n$ respectively. Let 
$S$ be an analytic subset of $M,$ we say that $S$ is \emph{thin} if $M\setminus S$ is dense in $M.$ 

\subsection{Jacobian bundles}~

  Let
$$f:M\rightarrow X$$
 be a
holomorphic mapping.
Given a holomorphic vector bundle $\pi:E\rightarrow X,$   the {\it
pull-back bundle}\index{pullback bundle}
$\tilde{\pi}:f^*E\rightarrow M$ is naturally defined up to an
isomorphism $\tilde f: f^*E_x\rightarrow E_{f(x)}$ for every $x\in M$
such that $f\circ\tilde{\pi}=\pi \circ\tilde{f}.$
Or equivalently,
 $$ f^*E=\Big\{(x,\xi)\in M\times E: \ f(x)=\pi(\xi)\Big\} $$
satisfying  that $\tilde{\pi}(x,\xi)=x$ and $\tilde{f}(x,\xi)=\xi.$  Let $H^0(X,E)$ stand for the space of all holomorphic sections of $E$ over $X$. For $s\in H^0(X,E)$,  a {\it lifted
section}\index{lifted section} $s_f\in H^0(M, f^*E)$ of $s$
by $f$ is  defined  by
\begin{equation}\label{lifted section}
 s_f(x)=\tilde{f}^{-1}\big(s\circ f(x)\big), \ \  ^\forall x\in M.
 \end{equation}
 If $s\not\equiv0$ with  $f(M)\not\subseteq s^{-1}(0),$  then  $s_f\not\equiv 0.$ So,  the \emph{pull-back
divisor} $f^*(s)$ of $(s)$ exists, and the multiplicity functions  of divisors $(s_f)$ and $f^*(s)$ satisfy
 $$\mu_{(s_f)}\geq \mu_{f^*(s)}.$$

The {\it Jacobian bundle}\index{Jacobian bundle} of $f$  is defined by
 $$K(f)=K_M\otimes f^*K_X^*,$$
where
$K_X^*$ is the dual of   canonical line bundle $K_X$.  A holomorphic section $F$ of $K(f)$
over $M$ is called a {\it Jacobian section of $f$}.  The section $F$ is said to be  {\it
effective}\index{effective} if $F^{-1}(0)$ is thin. The zero
divisor $(F)$   is called the {\it ramification divisor of $f$ with respect to $F.$}

Given  a Jacobian section $F$ of $f.$   Now we fix a  Hermitian inner product
on $K_X^*\oplus K_X,$ the inner product can
  pull  back to
 $f^*(K_X^*)\oplus f^*(K_X).$
In further, it induces  a $K_M$-valued Hermitian inner product
 $$\langle \cdot ,\cdot\rangle: \ K(f)\oplus f^*K_X\rightarrow K_M$$
 in a natural way.  Let $V$ be an open subset of $X$ such that $\tilde V:=f^{-1}(V)\not=\emptyset.$ The section $F$ defines a linear mapping (denoted  also by $F$)
 $$\ F: \  H^0(V,K_X)\rightarrow H^0(\tilde{V},K_M)$$
by $F[\Psi]=\langle F, \Psi_f\rangle$  for
$\Psi\in H^0(V,K_X)$.

Take an open subset $W$ of $X$ with $\tilde{W}:=f^{-1}(W)\not=\emptyset.$  Let  $(U; z_1,...,z_m),$   $(V; w_1,...,w_n)$
be holomorphic coordinate charts of $M, X$ respectively such that
$V\subseteq W$ and $f(U)\subseteq V.$  For $\Psi\in H^0(W,K_X),$ we write $\Psi|_V=\tilde {\Psi}dw$  and
$$F|_U=\tilde{F}dz\otimes d^*w_f,$$  where $\tilde{\Psi}, \tilde{F}$ are holomorphic functions on $V, U$ respectively,
and 
 $$dz=dz_1\wedge\cdots\wedge dz_m,\quad
 d^*w=\frac{\partial}{\partial w_1}\wedge\cdots\wedge \frac{\partial}{\partial
 w_n}.$$
By definition, it yields  that 
 $$F[\Psi]\big|_U=\tilde{\Psi}\circ f\cdot F[dw]=\tilde{\Psi}\circ f\cdot \tilde{F}dz.$$
 Let $A^{2n}(W)$ denote the set of smooth $2n$-forms on $W.$  We  
extend  $F$ linearly to
 $$F: \ A^{2n}(W)\rightarrow A^{2m}(\tilde{W})$$
in  a natural way: if $\Omega\in A^{2n}(W)$  is
 expressed  locally as
 $\Omega|_V=i_n\rho dw\wedge d\bar{w},$ where 
 $$i_p=\Big(\frac{\sqrt{-1}}{2\pi}\Big)^p(-1)^{\frac{p(p-1)}{2}} p!, \ \ ^\forall p\geq1,$$
 then 
$$ F[\Omega]\big|_U=i_m\rho\circ
 f\cdot F[dw]\wedge\overline{F[dw]}=i_m\rho\circ
 f\cdot |\tilde{F}|^2dz\wedge d\bar{z}.$$
Clearly,  we have
$$F[g_1\Omega_1+g_2\Omega_2]=g_1\circ f\cdot F[\Omega_1]+g_2\circ f\cdot F[\Omega_2]$$
for  $g_1,g_2\in A^0(W),$  $\Omega_1,\Omega_2\in A^{2n}(W).$
Let $\kappa$ be  a Hermitian metric on $K(f),$ then 
$\|F\|^2_\kappa=\tilde\kappa|\tilde{F}|^2$ with $\tilde\kappa|_U=\|dz\otimes
d^*w_f\|^2_\kappa.$ There is  a form $\Theta\in A^{2m}(\tilde{W})$
such that
\begin{equation*}
F[\Omega]=\|F\|^2_\kappa\Theta
\end{equation*}
with
$$\Theta|_U=i_m\rho\circ
 f\cdot\tilde\kappa^{-1}dz\wedge d\bar{z}.$$
Obviously, $\Theta>0$ if and only if $\Omega>0$.

In what follows, we  give an extension of Jacobian sections. 
  \begin{defi}\label{def}
Let $E$ be a holomorphic line bundle over $X.$ A holomorphic section $F_E$ of $f^*E\otimes K(f)$
over $M$ is called  a {\it Jacobian
section of $f$ with respect to $E$.}  We say that $F_E$ is  {\it
effective}\index{effective} if $F^{-1}_E(0)$  is thin. The zero
divisor $(F_E)$ is called the {\it ramification
divisor of $f$  with respect to $F_E$.}
  \end{defi}\label{def}
Let  $F_E$  be a Jacobian section of $f$ with respect to $E,$ we define   a   $f^*E\otimes K_M$-valued  interior product
 $$\langle \cdot,\cdot\rangle: \ \big(f^*E\otimes K(f)\big)\oplus f^*K_X\rightarrow f^*E\otimes K_M$$
induced  from  the natural    interior  product
on $K_X^*\oplus K_X.$
 Then $F_E$  defines  a linear mapping
 $$F_E: \ H^0(V, K_X)\rightarrow H^0(\tilde{V},f^*E\otimes K_M)$$
by $F_E[\Psi]=\langle F_E, \Psi_f\rangle$ for
$\Psi\in\Gamma(V, K_X)$. Let $\sigma$ be a local holomorphic frame of $E$ restricted to $V,$  and  write $F_E$ as
$$F_E|_U=\tilde{F}_E\sigma_f\otimes dz\otimes d^*w_f.$$  Let $\Psi\in H^0(W,K_X)$ with the  expression given  before,  then we have 
 $$F_E[\Psi]\big|_U=\tilde{\Psi}\circ f\cdot \sigma_f\otimes F_E[dw]=\tilde{\Psi}\circ f \cdot\tilde{F}_E\cdot\sigma_f\otimes dz.$$
Equip $E$ with  a Hermitian metric.  Similarly,  $F_E$  can  extend linearly to
 $$F_E: \ A^{2n}(W)\rightarrow A^{2m}(\tilde{W})$$
by  
\begin{equation}\label{qq}
 F_E[\Omega]\big|_U=i_m\rho\circ
 f\cdot \|\sigma_f\|^2\cdot|\tilde{F}_E|^2dz\wedge d\overline{z}
\end{equation}
for  $\Omega|_V=i_n\rho dw\wedge d\overline{w}\in A^{2n}(W).$
If $\kappa_E$ is  a Hermitian metric on  $f^*E\otimes K(f),$ then 
$\|F_E\|^2_{\kappa_E}=\tilde\kappa_E|\tilde{F}_E|^2$  with  $\tilde\kappa_E|_U=\|\sigma_f\otimes dz\otimes
d^*w_f\|^2.$ Also, there exists a form $\Theta\in A^{2m}(\tilde{W})$
such that
$$F_E[\Omega]=\|F_E\|^2_{\kappa_E}\Theta$$
with
$$\Theta|_U=i_m\rho\circ
 f\cdot\tilde\kappa_E^{-1}dz\wedge d\bar{z}.
$$
It is clear   that  $\Theta>0$ if and only if $\Omega>0$.

\subsection{Holomorphic fields} ~

 Let $E$  be a holomorphic line bundle over $X.$ Now we begin with  the notion of ``\emph{linearly almost-independent}" and ``\emph{pole order at most $E$}" for 
 vector fields.
  \begin{defi}\label{def}
 Let  $\mathfrak X_1,\cdots, \mathfrak X_k$  be meromorphic vector fields over $X.$
 We say that $\mathfrak X_1, \cdots, \mathfrak X_k$  are linearly almost-independent if 
 $\mathfrak X_1\wedge\cdots\wedge \mathfrak X_k\not\equiv0.$
We say that $\mathfrak X_1\wedge\cdots\wedge \mathfrak X_k$  is of   pole order at most $E$,  if  there exists a nonzero holomorphic section $t: X\rightarrow E$  such that
$t\otimes \mathfrak X_1\wedge\cdots\wedge \mathfrak X_k$
 is holomorphic.  For such $t,$ we have
$$
  \big(t\otimes \mathfrak X_1\wedge\cdots\wedge \mathfrak X_k\big)_f\in H^0\Big(M,f^*\big(E\otimes\bigwedge^{k}T_X\big)\Big).
$$
 \end{defi}
 Assume that $m<n$ and set
$q:=n-m.$ Apparently,  there exists a unique homomorphism
 $$\hat{f}: \ f^*\Big(E\otimes\bigwedge^{m}T^*_X\Big)\rightarrow
 f^*E\otimes\bigwedge^{m}T^*_M$$
 such that $\hat{f}(\xi_f)=f^*\xi$ for all $\xi\in
H^0\left(V,E\otimes\wedge^{m}T^*_X\right),$ where $V$ is  open  in $X$ with
$\tilde{V}:=f^{-1}(V)\not=\emptyset.$
Recall that  the interior
product
 $$\angle: \ K_X\oplus \Big(E\otimes \bigwedge^{q}T_X\Big)\rightarrow E\otimes \bigwedge^{m}
 T_X^*$$
is  defined by $\theta(\alpha\angle \xi)=(\xi\wedge\theta)(\alpha)\in E$ for every $\theta\in \bigwedge^{m}
 T_X,$ where $\alpha\in K_X$ and
 $\xi\in E\otimes \bigwedge^{q}T_N.$
 The interior product
  pulls back to an interior product
 $$\angle: \ f^*K_X\oplus f^*\Big(E\otimes\bigwedge^{q}T_X\Big)\rightarrow
 f^*\Big(E\otimes\bigwedge^{m}T^*_X\Big).$$
 
We call  $\varphi\in
H^0\left(M,f^*\left(E\otimes\bigwedge^{q}T_X\right)\right)$
 a {\it holomorphic field of
$f$ over $M$ of degree $q$ with respect to $E.$}  The section  $\varphi$ defines  a Jacobian
section $F_\varphi$ of $f$ with respect to $E$  by
 $$F_\varphi|_{\tilde{V}}=\hat{f}(\Psi_f\angle\varphi)\otimes
 \Psi_f^*,$$
 where $\Psi\in H^0(V,K_X)$ vanishes nowhere on $V,$  and $\Psi^*$ denotes
the dual  of $\Psi.$ Equivalently,  $F_\varphi$ is described by
 $F_\varphi[\Psi]= \hat{f}(\Psi_f\angle \varphi)$
  for every $\Psi\in H^0(V,K_X)$ with  $\tilde{V}\not=\emptyset$ for $V$ open in $X.$
 We say that $\varphi$ is 
  {\it effective}\index{effective} if $F_\varphi$ is
effective.

Let us  treat the existence of effective holomorphic fields of a holomorphic mapping $f: M\rightarrow X$ under the dimension condition $q:=n-m>0.$ Fix an integer $k$ such that $1\leq k\leq n,$ and 
 denote by $J_{1,k}^n$  be  the set of  all increasing injective mappings
\begin{equation*}
    \lambda: \ \mathbb{Z}[1,k]\longrightarrow \mathbb{Z}[1,n],
\end{equation*}
where $\mathbb{Z}[r,s]$ ($r\leq s$) denotes  the set of  integers $j$ with  $r\leq j\leq s$.
If $\lambda\in J_{1,k}^n,$ then $\lambda^{\bot}$ is uniquely defined such that $(\lambda, \lambda^{\bot})$ is a permutation of $\{1, \cdots,n\}.$ Clearly,
 $\bot: J_{1,k}^n\rightarrow  J_{1, n-k}^n$ is a bijective mapping.

 \begin{lemma}\label{lem0}  Assume $M$ is stein. Then a holomorphic field $\varphi$ of $f$ over $M$ of
 degree $q$ with respect to $E$ exists such that $\varphi$ is effective if and only if $f$ is differentiably non-degenerate.
\end{lemma}
\begin{proof}
 Let $S$  be the set of  all $x\in M$ such that
the rank of Jacobian  matrix of $f$ at $x$ is smaller than $m.$  Fix  $x_0\in M,$
we take local holomorphic coordinate
charts $(U;z_1,\cdots,z_m)$ of $x_0$ in $M,$ and $(V; w_1,\cdots,w_n)$ of
$f(x_0)$ in $X$ with $f(U)\subseteq V.$ We may assume that $E|_V\cong V\times\mathbb C.$
Set
$$d^*w_j=\frac{\partial}{\partial w_j}, \ \ j=1,\cdots, n.$$
For $\nu\in J^n_{1,m},$  we write
$f^*dw_\nu=A_\nu dz,$
where
 $dw_\nu=dw_{\nu(1)}\wedge\cdots\wedge d_{\nu(m)}$ and $dz=dz_1\wedge\cdots dz_m.$
By  definition of ranks,  it yields  that
$$S\cap U=\bigcap_{\nu\in J^n_{1,m}}A^{-1}_\nu(0).$$
\ \ \ \ \emph{Necessarity.} Assume  that $\varphi$ is an effective field of $f$ over $M$ of
 degree $q$ with respect to $E,$ we  prove that $f$ is differentiably non-degenerate.
 Let $\sigma$ be a local holomorphic
 frame of $E$ restricted to $V,$  we write
$$
\varphi|_U=\sum_{\lambda\in J^n_{1,q}}\varphi_\lambda \sigma_f\otimes d^*w_{\lambda f},
$$
where
$d^*w_{\lambda}=d^*w_{\lambda(1)}\wedge\cdots\wedge d^*w_{\lambda(q)}.$
 A simple computation shows that
$$dw_f\angle\varphi|_U=\sum_{\lambda\in J^n_{1,q}}{\rm sign}(\lambda^\bot, \lambda)
\varphi_\lambda \sigma_f\otimes dw_{\lambda^\bot f}.$$
Hence, we obtain
 $$F_\varphi[dw]\big|_U=
\hat{f}(dw_f\angle\varphi)=\sum_{\lambda\in J^n_{1,q}}{\rm sign}(\lambda^\bot, \lambda)
\varphi_\lambda A_{\lambda^\bot}\sigma_f\otimes dz,
$$
which implies that
$$(F_\varphi)\cap U=\Bigg(\sum_{\lambda\in J^n_{1,q}}{\rm sign}(\lambda^\bot, \lambda)
\varphi_\lambda A_{\lambda^\bot}\Bigg)\supseteq
S\cap U.$$
Since $\varphi$ is effective, then $S$ is thin. Thus, $f$ is differentiably non-degenerate.

\emph{Sufficiency.} Assume  that $f$ is differentiably non-degenerate,  we show that there exists an effective field of $f$ over $M$ of
 degree $q$ with respect to $E.$ 
For 
$x_0\in M\setminus S,$    there exists    $\iota\in J_{1,m}^n$  satisfying
that $A_\iota(x_0)\not=0$.
Since $M$ is stein,   then  by Corollary 5.6.3   in \cite{Ho} we note that  there exists   $\tau\in H^0(M, f^*E)$
and $s_j\in H^0(M, f^*T_X)$  such that
$$\tau(x_0)\not=0; \ \ \ s_j(x_0)=d^*w_{jf}(x_0), \ \ j=1,\cdots,n.$$
Set $s_{\iota}=s_{\iota(1)}\wedge\cdots\wedge s_{\iota(m)},$  then 
$$\tau\otimes s_{\iota^{\bot}}\in H^0\Big(M,f^*E\otimes\bigwedge^{q}{f^*
T_X}\Big).$$
Now define $\varphi:=\tau\otimes s_{\iota^{\bot}},$
 which is a holomorphic
field of $f$ over $M$ of degree $q$ with respect to $E.$ Next, we prove  that $\varphi$ is effective.  Write
$$\varphi|_U=\sum_{\lambda\in J^n_{1,q}}\varphi_\lambda \tau\otimes (d^*w_{\lambda})_f,$$
which is satisfied with 
$$\varphi_\lambda(x_0)=1, \ \ \lambda=\iota^{\bot}; \ \ \ \varphi_\lambda(x_0)=0, \ \ \lambda\not=\iota^{\bot}.$$
So, it yields  that 
$$F_\varphi[dw]=\sum_{\lambda\in J^n_{1,q}}{\rm sign}(\lambda^\bot, \lambda)
\varphi_\lambda A_{\lambda^\bot}\tau\otimes dz
$$
with
$$F_\varphi[dw](x_0)={\rm sign}(\iota, \iota^\bot)A_{\iota}(x_0)\tau(x_0)\otimes dz(x_0)\not=0.
$$
Notice that
$$(F_\varphi[dw])\cap U=(F_\varphi)\cap U,$$
hence $x_0$ is not a zero of $F_\varphi,$ which implies that $ {\rm Supp}(F_\varphi)\subseteq S$ is thin since $S$ is thin. Therefore, $\varphi$ is  effective. We conclude the proof.
 \end{proof}

\begin{theorem}
Assume   $f$ is differentiablely non-degenerate.  Let $\mathfrak X_1,\cdots, \mathfrak X_n$ be linearly almost-independent
 meromorphic vector fields  over $X,$ and let $\mathfrak X_1\wedge\cdots\wedge \mathfrak X_n$ be of pole order at most $E,$ i.e.,  there is
 a nonzero holomorphic section $t: X\rightarrow E$ such that $t\otimes \mathscr X$ is holomorphic with 
$\mathscr X:=\mathfrak X_1\wedge\cdots \wedge \mathfrak X_n.$
 If $f(M)\nsubseteq {\rm Supp}(t\otimes \mathscr X),$
then there is
$\lambda\in J_{1,q}^n$ such that
 $(t\otimes \mathfrak X_{\lambda(1)}\wedge\cdots\wedge
 \mathfrak X_{\lambda(q)})_f$
is an effective holomorphic field of $f$ over $M$ of degree $q$ with respect to $E$.
\end{theorem}
\begin{proof}
The conditions imply that  the lifted section
$$t_f\otimes\mathscr X_{f}\in H^0(M,f^*(E\otimes K_X^*))$$
   exists with
$t_f\otimes\mathscr X_{f}\not\equiv 0$.  Let $S$  be the set of  all $x\in M$ such that
the rank of Jacobian  matrix of $f$ at $x$ is smaller than $m.$
 Note  that   $S, (\mathscr X_t), (t_f\otimes\mathscr X_{f})$ are  thin.
Take $x_0\in
M\setminus S$ such that $f(x_0)\not\in {\rm Supp}(t\otimes\mathscr X),$
then there exists local holomorphic coordinate
charts $(U;z_1,\cdots,z_m)$ of $x_0$ and $(V; w_1,\cdots,w_n)$ of
$f(x_0)$ with $f(U)\subseteq V,$ such that  $E|_V\cong V\times\mathbb C$ and $t\otimes \mathfrak X_1\wedge \cdots\wedge\otimes \mathfrak X_n\not=0$ on $V.$ 
Write $t=\tilde t\sigma,$ where  $\sigma$ is a local holomorphic
 frame of $E$ restricted to $V.$ One can factorize $\tilde t$ as $\tilde t=\tilde t_1\cdots\tilde t_n$ such that $\tilde t_1\mathfrak X_1,\cdots,\tilde t_n\mathfrak X_n$ are holomorphic  and have no zeros on $V,$ and hence 
form a local holomorphic frame of $T_X$ on $V.$ 
Let $\omega_1,\cdots, \omega_n$ denote   
  the dual frame relative to  $ \tilde t_1\mathfrak X_1,\cdots, \tilde t_n\mathfrak X_n$ restricted to $V.$
For $\nu\in J_{1,m}^{n},$  we write
\begin{equation*}
f^*\omega_\nu|_U=B_\nu dz,
\end{equation*}
 where
$\omega_\nu=\omega_{\nu(1)}\wedge\cdots\wedge
\omega_{\nu(m)}.$
 Then there exists  $\iota\in J_{1,m}^n$  such
that $B_\iota(x_0)\not=0$. Take  $\lambda=\iota^\bot\in J_{1,q}^n$ and 
  define
 $$\varphi:=\big(t\otimes \mathfrak X_{\lambda(1)}\wedge\cdots\wedge
 \mathfrak X_{\lambda(q)}\big)_f.$$
 Clearly, 
$\varphi$ is a holomorphic field of $f$ over $M$ of degree $q$ with respect to $E.$
Next, we show  that $\varphi$ is effective.  Set  $$\Psi=\omega_1\wedge\cdots\wedge \omega_n.$$
We have  
 $$\Psi_f\angle\varphi={\rm sign}(\iota,\iota^\bot) t_f\otimes \omega_{\iota
 f}$$
with $\omega_{\iota f}=(\omega_{\iota (1)} \wedge\cdots\wedge\omega_{\iota (m)})_f,$ which yields that
\begin{equation*}
F_\varphi[\Psi]\big|_U= \hat{f}(\Psi_f\angle \varphi)={\rm
sign}(\iota,\iota^\bot) B_{\iota}t_{f}\otimes dz.
\end{equation*}
So, $x_0\in M\setminus{\rm Supp}(F_\varphi)$ due to $t_f|_U\not=0$ and
$$(F_\varphi[\Psi])\cap U=(F_\varphi)\cap U.$$
That is to say,  ${\rm Supp}(F_\varphi)\subseteq S\cup\mathscr (t_f\otimes \mathscr X_{f})$  is thin. Hence,
$\varphi$ is effective.
\end{proof}

\begin{theorem}\label{thm3}
Let   $M=\mathbb{C}^m.$ Set
 $$\mathfrak Z_j=f_*\Big(\frac{\partial}{\partial z_j}\Big), \ \
 j=1,\cdots,m.$$
Then a holomorphic field $\varphi$ of $f$ over
$\mathbb{C}^m$ of degree $q$ with respect to $E$ is effective  if and only if
 $$\mathscr Z:=\mathfrak Z_{1f}\wedge\cdots\wedge \mathfrak Z_{mf}\wedge \varphi\not\equiv 0.$$
Moreover, we have $(F_\varphi)=(\mathscr Z)$.
\end{theorem}

\begin{proof}
  Take a
holomorphic coordinate chart  $(V; w_1,\cdots,w_n)$
of $X.$ We may assume that $E|_V\cong V\times \mathbb C.$
 Write
 $$f_j=w_j\circ f, \ \
 j=1,\cdots,n; \ \ \  \mathfrak Z_i=\sum_{k=1}^n\frac{\partial f_k}{\partial
 z_i}d^*w_k,\ \  i=1,\cdots,m,$$
 where $d^*w_k=\partial/\partial w_k$ for $1\leq k\leq m.$
A simple computation gives that
 $$\mathfrak Z_{1f}\wedge\cdots\wedge \mathfrak Z_{mf}=\sum_{\nu\in
 J_{1,m}^{n}}A_\nu d^*w_{\nu f},$$
where 
 $$A_\nu=\det\Big(\frac{\partial f_{\nu(j)}}{\partial
 z_i}\Big), \ \ \ d^*w_{\nu}=d^*w_{\nu(1)}\wedge\cdots\wedge d^*w_{\nu(m)}.$$
 Let $\sigma$ be a local holomorphic
 frame of $E$ on  $V$ and write  
$$
\varphi|_{f^{-1}(V)}=\sum_{\lambda\in J^n_{1,q}}\varphi_\lambda \sigma_f\otimes d^*w_{\lambda f},
$$
 where  
$d^*w_{\lambda}=d^*w_{\lambda(1)}\wedge\cdots\wedge d^*w_{\lambda(q)}.$ Hence, we obtain
$$\mathscr Z|_{f^{-1}(V)}=\sum_{\lambda\in J^n_{1,q}}{\rm sign}(\lambda^{\bot}, \lambda)
\varphi_\lambda A_{\lambda^{\bot}}\sigma_f\otimes d^*w_f.$$
On the other hand, we have
 $$F_\varphi[dw]\big|_{f^{-1}(V)}=
 \sum_{\lambda\in J_{1,q}^{n}}{\rm sign}(\lambda^\perp,\lambda)\varphi_\lambda
 A_{\lambda^\perp}\sigma_f\otimes dz,$$
 where $dz=dz_1\wedge\cdots\wedge dz_m.$
So,  we conclude that  
 $$(F_\varphi)\cap f^{-1}(V)=(F_\varphi[dw])\cap f^{-1}(V)=(\mathscr Z)\cap f^{-1}(V).$$
 The proof is completed. 
\end{proof}
\begin{theorem}\label{lem2} Assume  $f$ is differentiably non-degenerate.  Let  $\varphi$ be a holomorphic field
of $f$ over $M$ of degree $q$ with respect to $E.$ Equip  $M, X, E$ with Hermitian metrics $\alpha,\omega, h$
 respectively. Define
 a non-negative function $g$ 
by
 $$F_\varphi[\omega^n]=g^2f^*\omega^m.$$
Then $g\leq \|\varphi\|,$  where $\|\cdot\|$ is induced from $\omega, h.$
\end{theorem}
\begin{proof}

For any  $x_0\in M,$
we can  take a  local holomorphic frame
$(U; \theta_1,\cdots,\theta_m)$ of $T^*_M$ around  $x_0$ and a local holomorphic frame
 $(V; \psi_1,\cdots,\psi_n)$  of  $T^*_X$ around  $f(x_0)$ with   $f(U)\subseteq V,$ such that
 $$\alpha|_{x_0}=\frac{\sqrt{-1}}{2\pi}
 \sum_{j=1}^m\theta_j\wedge
 \bar{\theta}_j, \ \ \omega|_{f(x_0)}=\frac{\sqrt{-1}}{2\pi}\sum_{j=1}^n\psi_j\wedge
 \bar{\psi}_j.$$
In addition, we may assume that $E|_V\cong V\times \mathbb C.$ Set
$$ f^*\omega^m=u\alpha^m; \ \ \   f^*\psi_\nu=A_{\nu}\theta_1\wedge\cdots\wedge\theta_m,\ \  ^\forall \nu\in J_{1,m}^{n}
$$
with 
 $\psi_\nu=\psi_{\nu(1)}\wedge
\psi_{\nu(2)}\wedge\cdots
 \wedge \psi_{\nu(m)}.$
It is trivial to confirm that
\begin{equation}\label{drt}
 u(x_0)=\sum_{\nu\in J_{1,m}^{n}}|A_\nu(x_0)|^2\not=0.
 \end{equation}
   Let  $\sigma$ be a local holomorphic frame of $E$ and  write  
   \begin{equation}\label{123aw}
  \varphi|_U=\sum_{\lambda\in J_{1,q}^{n}}\varphi_\lambda\sigma_f\otimes
 \psi^*_{\lambda f}
\end{equation}
 with
 $\psi^*_\lambda=\psi^*_{\lambda(1)}\wedge\cdots
 \wedge \psi^*_{\lambda(q)},$
 where  $\psi^*_1,\cdots,\psi^*_n$ is  the dual frame relative to
$\psi_1,\cdots,\psi_n.$
Set 
$\Psi=\psi_1\wedge\cdots\wedge\psi_n.$ Then
 $$\Psi_f\angle\varphi=\sum_{\lambda\in J_{1,q}^{n}}{\rm sign}(\lambda^\perp,\lambda)\varphi_\lambda
\sigma_f\otimes\psi_{\lambda^\perp f}.$$
Consequently,
$$ F_\varphi[\Psi]\big|_U=\hat{f}(\Psi_f\angle\varphi)=
 \sum_{\lambda\in J_{1,q}^{n}}{\rm sign}(\lambda^\perp,\lambda)\varphi_\lambda
 A_{\lambda^\perp}\sigma_f\otimes\theta_1\wedge\cdots\wedge\theta_m.
$$
This leads to
 $$g^2(x_0)u(x_0)=\Bigg| \sum_{\lambda\in J_{1,q}^{n}}{\rm sign}(\lambda^\perp,\lambda)\varphi_\lambda(x_0)
 A_{\lambda^\perp}(x_0)\Bigg|^2\|\sigma_f(x_0)\|^2.$$
Combine  (\ref{drt}) and (\ref{123aw}) with Schwarz's inequality,
 $$\Bigg| \sum_{\lambda\in J_{1,q}^{n}}{\rm sign}(\lambda^\perp,\lambda)\varphi_\lambda(x_0)
 A_{\lambda^\perp}(x_0)\Bigg|^2\|\sigma_f(x_0)\|^2\leq u(x_0)\|\varphi(x_0)\|^2$$
which follows that  $g(x_0)\leq \|\varphi(x_0)\|.$ This proves the theorem.
\end{proof}

\section{Second Main Theorem type inequalities}

\subsection{Preliminaries}~

For $z=(z_1,\cdots,z_m)\in\mathbb C^m,$  set $\|z\|^2=|z_1|^2+\cdots+|z_m|^2$ and 
$$\alpha=dd^c\|z\|^2,\ \ \ \gamma=d^c\log\|z\|^2\wedge \left(dd^c\log\|z\|^2\right)^{m-1}$$
with $d^c=\frac{\sqrt{-1}}{4\pi}(\bar\partial-\partial)$ and $dd^c=\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}.$
Let $$f:\mathbb C^m\rightarrow X$$  be a holomorphic mapping, where $X$  is a  smooth complex projective variety.
Given  an ample  line bundle  $L\rightarrow X$
  which carries  a Hermitian metric $h.$
 The \emph{characteristic function} of $f$ with respect to $L$ is defined by
 $$ T_f(r,L)=\int_1^r\frac{dt}{t^{2m-1}}\int_{\|z\|<t}f^*c_1(L,h)\wedge\alpha^{m-1},$$
 up to a bounded term for different metrics $h$ on $L.$   Since a holomorphic line bundle  on $X$
 can be written as the difference of two  ample  line bundles, then   definition of $T_{f}(r,L)$   extends to
 an arbitrary holomorphic line bundle.  Let $D$ be a  divisor on $X,$
   define   \emph{proximity function} of $f$ with respect to $D$ by
$$m_f(r,D)=\int_{\|z\|=r}\log\frac{1}{\|s_D\circ f\|}\gamma,$$
where  $s_D$ is the canonical section associated to $D.$  Write $s_D=\tilde{s}_De$ locally, where $e$ is a local holomorphic frame of $([D],h).$  We define   \emph{counting function}
 of $f$ with respect to $D$ by
$$N_f(r,D)=\int_1^r\frac{dt}{t^{2m-1}}\int_{\|z\|=t}dd^c\log|\tilde{s}_D\circ f|^2\wedge\alpha^{m-1}.$$
Assume  that $f(\mathbb C^m)\not\subseteq D,$ then the First Main Theorem (cf. \cite{ru}) says that
$$T_f(r,[D])=m_f(r,D)+N_f(r,D)+O(1).$$
\begin{lemma}[Borel Lemma, \cite{No}] Let $\phi$ be a monotone increasing function on $[0,\infty)$ such that $\phi(r_0)>1$ for some $r_0\geq0.$
Then for any $\delta>0,$ there exists a set $E_\delta\subset[0,\infty)$ of finite Lebegue measure such that
$$\phi'(r)\leq \phi(r)\log^{1+\delta}\phi(r)$$
holds for $r>0$ outside $E_\delta.$
\end{lemma}
\begin{proof} Since $\phi$ is monotone increasing,
$\phi'(r)$ exists for almost all $r\geq0.$
Set
$$S=\left\{r\geq0: \phi'(r)>\phi(r)\log^{1+\delta}\phi(r)\right\}.$$
Then
 \begin{eqnarray*}
\int_Sdr &\leq&\int_0^{r_0}dr+\int_{S\setminus[0,r_0]}dr
\leq r_0+\int_{r_0}^\infty\frac{\phi'(r)}{\phi(r)\log^{1+\delta}\phi(r)}dr <\infty.
 \end{eqnarray*}
\end{proof}
Let $\kappa$ be a non-negative, locally integrable real-valued function on $\mathbb C^m.$ Set
$$T_\kappa(r)=\int_1^r\frac{dt}{t^{2m-1}}\int_{\|z\|<t}\kappa\alpha^m.$$
Then $T_\kappa(r)$ is a monotone increasing function on $[1,\infty).$
\begin{lemma}[Calculus Lemma]\label{cal} Let $\kappa$ be a non-negative, locally integrable  real-valued function  on $\mathbb C^m$ such that  
$r^{2m-1}T'_\kappa(r)$  is monotone increasing in $r$ and  $r_0^{2m-1}T'_\kappa(r_0)>1$ for some $r_0\geq1.$
Then
$$\log^+\int_{\|z\|=r}\kappa\gamma\leq_{\rm exc}
O\Big(\log^+ T_\kappa(r)+\log^+\log r\Big).$$
\end{lemma}
\begin{proof} Denote by $dV$ the  Euclidean volume measure on $\mathbb C^m,$ and  by
$d\sigma_r$  the   volume measure (induced by $dV$) on $\{\|z\|=r\}$ for $r>0.$ A well-known fact says that
$$dV=\frac{\pi^m}{m!}\alpha^m, \ \ \  d\sigma_r=\omega_{2m-1}r^{2m-1}\gamma,$$
where $\omega_{2m-1}$ is the Euclidean volume of  unit sphere in $\mathbb R^{2m}.$ Therefore,
\begin{eqnarray*}
\frac{d}{dr}T_\kappa(r)&=&\frac{1}{r^{2m-1}}\int_{\|z\|<r}\kappa\alpha^m
= \frac{m!}{\pi^mr^{2m-1}}\int_{\|z\|<r}\kappa dV \\
&=& \frac{m!}{\pi^mr^{2m-1}}\int_0^rdt\int_{\|z\|=t}\kappa d\sigma_t \\
&=& \frac{m!\omega_{2m-1}}{\pi^mr^{2m-1}}\int_0^rdt\int_{\|z\|=t}\kappa \gamma,
 \end{eqnarray*}
which results in
$$\int_{\|z\|=r}\kappa \gamma=\frac{\pi^m}{m!\omega_{2m-1}r^{2m-1}}\frac{d}{dr}\Big(r^{2m-1}\frac{d}{dr}T_\kappa(r)\Big).$$
Indeed, we can use Borel's lemma to get
\begin{eqnarray*}
\frac{d}{dr}T_\kappa(r)&\leq_{\rm exc}& T_\kappa(r)\log^{1+\delta}T_\kappa(r),  \\
\frac{d}{dr}\Big(r^{2m-1}\frac{d}{dr}T_\kappa(r)\Big)&\leq_{\rm exc}&
\Big(r^{2m-1}\frac{d}{dr}T_\kappa(r)\Big)
\log^{1+\delta}\Big(r^{2m-1}\frac{d}{dr}T_\kappa(r) \Big).
 \end{eqnarray*}
Put  the above together, we conclude   that
$$\log^+\int_{\|z\|=r}\kappa\gamma\leq_{\rm exc}
O\Big(\log^+ T_\kappa(r)+\log^+\log r\Big).$$
\end{proof}
\begin{lemma}[Green-Jensen formula, \cite{No}]\label{GJ} Let $u\not\equiv-\infty$ be a plurisubharmonic function on $\mathbb C^m.$ Then for  any $0<s<r,$
$$\int_{\|z\|=r}u\gamma-\int_{\|z\|=s}u\gamma=2\int_s^r\frac{dt}{t^{2m-1}}\int_{\|z\|<t}dd^cu\wedge\alpha^{m-1}$$
holds in the sense of currents.
\end{lemma}

\subsection{Second Main Theorem type inequalities}~

Let $$f:\mathbb C^m\rightarrow X$$
 be a holomorphic mapping into a smooth complex projective variety $X$ with complex dimension $n>m.$
Let $E$ be  a holomorphic line bundle over $X.$  For $\varphi,$  an effective holomorphic field
of $f$ over $M$ of degree $q$ with respect to $E,$   
   the  {\it ramification
term} $N_{f, {\rm Ram}}(r)$ of $f$ with respect to $F_\varphi$ is defined by 
$$N_{f, {\rm Ram}}(r)=N(r, (F_\varphi)).
$$

\begin{theorem}\label{mainthm}
Let $E$ be a holomorphic line bundle over $X,$ and let $L$ be an ample line bundle over $X.$
  Let $f:\mathbb C^m\rightarrow X$ be a  holomorphic mapping.   Set
 $$\mathfrak Z_j=f_*\Big(\frac{\partial}{\partial z_j}\Big), \ \
 j=1,\cdots,m.$$
Let  $\mathfrak X_1,$ $\cdots, \mathfrak X_{n-m}$ be linearly almost-independent meromorphic vector fields over $X,$ and let $\mathfrak X_1\wedge \cdots\wedge \mathfrak X_{n-m}$ be
of pole order at most $E,$ i.e.,  there exists a nonzero holomorphic section
$t: X\rightarrow E$ such that $t\otimes \mathfrak X_1\wedge\cdots \wedge \mathfrak X_{n-m}$ is holomorphic.
Assume  that 
  $\mathfrak Z_{1f}\wedge\cdots\wedge \mathfrak Z_{mf}\wedge \varphi\not\equiv 0$
with  $\varphi=(t\otimes \mathfrak X_1\wedge\cdots \wedge \mathfrak X_{n-m})_f.$
If $f(\mathbb C^m)\nsubseteq F^{-1}_\varphi(0)$, then
\begin{equation*}
T_f(r,K_X)+ N_{f, {\rm Ram}}(r)\leq_{\rm exc} T_f(r,E)+ O\Big(\log T_f(r,L)+\log^+\log r\Big).
\end{equation*}
\end{theorem}

\begin{proof}  It is known  from Theorem \ref{thm3} that $\varphi$ is an effective holomorphic field
of $f$ over $M$ of degree $q$ with respect to $E.$  By Lemma \ref{lem0}, $f$ is differentiably non-degenerate. Since $L$ is ample, then there exists a Hermitian metric $h$ on $L$ such that 
$\omega:=c_1(L,h)>0.$
Define a non-negative function $\xi$ of  $C^\infty$-class on $\mathbb C^m\setminus F^{-1}_\varphi(0)$  by
\begin{equation}\label{xi}
    F_\varphi[\omega^n]=\xi\alpha^m.
\end{equation}
Let $h_E$ be a Hermitian metric on $E.$  Observe  (\ref{qq}),  one obtains
$${\rm{Ric}}( F_\varphi[\omega^n])=f^*{\rm{Ric}}(\omega^n)+f^*c_1(E, h_E).$$
This  yields  from Poincar\'e-Lelong formula that
\begin{equation*}
    dd^c\log \xi= f^*c_1(K_N,h_K) -f^*c_1(E, h_E)+(F_\varphi)
\end{equation*}
in the sense of currents, here $h_K$ is the Hermitian metric on $K_X$ induced  by $\omega.$
Integrating both sides of  the current  equation above,  we get
\begin{eqnarray*}
&& \int_1^r\frac{dt}{t^{2m-1}}\int_{\|z\|<t}dd^c\log \xi\wedge\alpha^{m-1} \\  &=&
 T_f(r,K_X) -T_f(r,E)+N_{f, {\rm Ram}}(r)+O(1).
\end{eqnarray*}
Apply Green-Jensen formula again, we arrive at
\begin{eqnarray}\label{5q}
\frac{1}{2}\int_{\|z\|=r}\log \xi\gamma   &=& T_f(r,K_X)-T_f(r,E)+N_{f, {\rm Ram}}(r)+O(1).
\end{eqnarray}
On the other hand,
set
 $$F_\varphi[\omega^n]=g^2f^*\omega^m.$$
Since $V$ is compact,  by  definition of $\varphi$ and Theorem \ref{lem2},  we have
 $g\leq \|\varphi\|<c$ for a positive constant $c.$
 Combined with (\ref{xi}), we obtain
 $$\xi=g^2\frac{f^*\omega^m}{\alpha^m}$$
 which yields  that
  $$\int_{\|z\|=r}\log\xi\gamma=\int_{\|z\|=r}\log g^2\gamma+\int_{\|z\|=r}\log\frac{f^*\omega^m}{\alpha^m}\gamma.$$
 The boundedness of $g$ means that
 $$\int_{\|z\|=r}\log g^2\gamma\leq O(1).$$
Now we claim: 
 \begin{equation}\label{claim}
 \int_{\|z\|=r}\log\frac{f^*\omega^m}{\alpha^m}\gamma\leq_{\rm exc} O\big(\log T_f(r,L)+\log^+\log r\big).
 \end{equation}
By this  with (\ref{5q}),  we  can confirm the theorem.
Set 
\begin{equation*}
  f^*\omega\wedge\alpha^{m-1}=\varrho\alpha^m.
\end{equation*}
For any  $z_0\in \mathbb C^m,$ one may pick  a holomorphic coordinate system $w_1,\cdots,w_n$ near $f(z_0)$ such that
\begin{equation*}
 \omega|_{f(z_0)}=\frac{\sqrt{-1}}{2\pi}\sum_{j=1}^n dw_j\wedge d\bar{w}_j.
\end{equation*}
Set $f_j=w_j\circ f$ for $1\leq j\leq n,$   then
\begin{equation*}
    f^*\omega\wedge\alpha^{m-1}|_{z_0}=\frac{(m-1)!}{2}\sum_{j=1}^n\big\|\nabla f_j\big\|^2\alpha^m,
\end{equation*}
where $\nabla$ is the gradient operator on $\mathbb C^m.$
That is,
$$    \varrho|_{z_0}=\frac{(m-1)!}{2}\sum_{j=1}^n\big\|\nabla f_j\big\|^2.
$$Thus, we have 
$$\frac{f^*\omega^m}{\alpha^m}\big|_{z_0}\leq B\varrho^m$$
holds for a  positive number $B.$
By the arbitrariness of $z_0,$ we receive
$$\log\frac{f^*\omega^m}{\alpha^m}\leq m\log \varrho+O(1)$$
 on $\mathbb C^m.$ Indeed, by Calculus Lemma (Lemma \ref{cal})
\begin{eqnarray*}
\int_{\|z\|=r}\log\varrho\gamma
&\leq& \log\int_{\|z\|=r}\varrho\gamma \\
&=& \log \int_{\|z\|=r}\frac{f^*\omega\wedge\alpha^{m-1}}{\alpha^m}\gamma
\\
&\leq_{\rm exc}& O\Big(\log\int_1^r\frac{dt}{t^{2m-1}}\int_{\|z\|<t}f^*\omega\wedge\alpha^{m-1}+\log^+\log r\Big) \\
   &=& O\big(\log T_f(r,L)+\log^+\log r\big).
\end{eqnarray*}
This conforms the claim (\ref{claim}).  We conclude the proof.
\end{proof}

\begin{lemma}[\cite{HY}]\label{SLang87lem5.2}
 Let
$H$ be a very ample line  bundle over $X$ and let $L$ be a pseudo
ample   line bundle over $X$. Then $$\limsup_{j\to +\infty}
\frac{1}{j^n}\dim H^0(X, L^j\otimes H^*)>0.$$
\end{lemma}

\begin{cor}\label{Cor1}
Let $E$ be a holomorphic line bundle over $X.$
 Set
 $$Z_j=f_*\Big(\frac{\partial}{\partial z_j}\Big), \ \
 j=1,\cdots,m.$$
Let $\mathfrak X_1,$ $\cdots, \mathfrak X_{n-m}$  be  linearly almost-independent meromorphic vector fields over $X,$
and let $\mathfrak X_1\wedge \cdots\wedge \mathfrak X_{n-m}$ be
of  pole order at most $E.$ If $E^*\otimes K_X$ is pseudo ample,  then every holomorphic mapping
 $f:\mathbb C^m\rightarrow X$  is algebraically degenerate or  satisfied with the equation
  $$\mathfrak Z_{1f}\wedge\cdots\wedge \mathfrak Z_{mf}\wedge (\mathfrak X_{1}\wedge\cdots\wedge \mathfrak X_{n-m})_f= 0.$$
 \end{cor}
 
\begin{proof} 
Since $N$ is projective algebraic, there exists a very ample line
bundle $H$ over $X$. By Lemma~\ref{SLang87lem5.2}, for $j$ large
there exists a non-trivial holomorphic section $s$ of $L^j\otimes
H^*$, where $L=E^*\otimes K_X$.  Since $X$ is compact, a Hermitian metric $\kappa$ on $L^j\otimes
H^*$ exists such that $\|s\|_\kappa\leq 1.$
Assume   $f(\mathbb C^m)\not\subseteq {\rm Supp}((s)),$ otherwise, the curve $f$ is algebraically degenerate.  By  First Main Theorem 
$$
N_f(r,(s))  \leq T_f(r,L^j\otimes H^*)+O(1)  =jT_f(r,L)-T_f(r,H)+O(1)
$$
which implies that 
\begin{equation}
 T_f(r,E^*\otimes K_X)\geq \frac{1}{j}T_f(r,H) +O(1),
\end{equation}
which contradicts with  
Theorem \ref{mainthm}.  
\end{proof}

 \begin{theorem}\label{ddc}
Let $X$ be a  $2$-dimensional smooth complex projective variety.  
Let $\mathfrak X$ be a nonzero meromorphic vector field of  pole order at most $E$  over $X.$  If $E^*\otimes K_X$ is pseudo ample, 
 then $X$  contains  no  algebraically non-degenerate holomorphic curves. 
\end{theorem}
\begin{proof} By the assumption, there is a nonzero  holomorphic section
$t: X\rightarrow E$ such that $t\otimes \mathfrak X$ is holomorphic. 
Note from Corllary \ref{Cor1} that 
$f$ is algebraically degenerate or  satisfied with the equation
\begin{equation}\label{s123}
\mathfrak Z_f\wedge \mathfrak X_f=0 \Longleftrightarrow \mathfrak Z_f\wedge (t\otimes\mathfrak X)_f=0.
\end{equation}
Let $\mathscr U$ be a finite open covering of $X$  such that every  element of $\mathscr U$ carries a holomorphic coordinate system. Take an element  $V\in\mathscr U$ with a  holomorphic coordinate system $(w_1,w_2)$ on $V.$ Locally, write 
$$t\otimes\mathfrak X|_V=A_1(w_1,w_2)\frac{\partial}{\partial w_{1}}+A_2(w_1,w_2)\frac{\partial}{\partial w_{2}},$$
 where $A_1,A_2$ are holomorphic in $w_1,w_2.$ If $f(\mathbb C)\cap V\not=\emptyset,$
then
$$\mathfrak Z|_{f(\mathbb C)\cap V}=f'_1\frac{\partial}{\partial w_{1}}+
f'_2\frac{\partial}{\partial w_{2}}$$
with $f_1=w_1\circ f$ and $f_2=w_2\circ f.$
Substitute the expressions of $t\otimes\mathfrak X$ and $\mathfrak Z$ into  (\ref{s123}),  it follows that
$$A_1(f_1,f_2)df_2+A_2(f_1, f_2)df_1=0.$$
Consider  the ODE
$$A_1(w_1,w_2)dw_2+A_2(w_1, w_2)dw_1=0.$$
Apply the integrating factor method in theory of ODEs,   there exists  $\Phi$ which  is holomorphic in $w_1,w_2,$ such that the equation is solved generally by
$$\Phi(w_1,w_2)=C,$$
where  $C$ is an arbitrary constant.  Thus,  we get
$$\Phi(f_1,f_2)=C_0$$
for a constant $C_0$ determined by $f.$
  This means  that $f|_{f^{-1}(V)}$ is algebraically degenerate. Since $\mathscr U$ is finite, then $f$ is algebraically degenerate.
\end{proof}

\subsection{Final notes}~

Let $X$ be a $n$-dimensional smooth complex projective  variety. 
Let $f:\mathbb{C}\rightarrow X$ be a nonconstant holomorphic
curve, the natural lifted curve  $f':\mathbb C\rightarrow T_X$ is the
derivative of $f.$  Take a local holomorphic coordinates $w=(w_1,...,w_n)$ of $X$ near
$f(z).$  For convenience,  we write  $f=(f_1,...,f_n),$  where $f_\alpha=w_\alpha\circ f.$  
Set $\partial_z=\partial /\partial
 z,$  $f'$ can be expressed as
 $$f'=\sum_{\alpha=1}^nf'_\alpha\partial_\alpha, \ \ f'_\alpha=\partial_zf_\alpha.$$  
\ \ \  \ Let $\mathscr D$ be a meromorphic connection on $T_X$ with
pole order at most $E,$ i.e., $t\otimes\mathscr D$ is holomorphic for a nonzero $t\in
H^0(X,E),$ where  $E$ is a holomorphic line bundle over $X.$ Equip a Hermitian metric on $E.$ For a
positive integer $k,$ define
 $$f^{(k+1)}=(t\otimes\mathscr D_{f'})^kf': \ \mathbb{C}\rightarrow E^k\otimes 
 T_X,$$
where $\mathscr D_{f'}$ is the covariant derivative with respect to the
connection $\mathscr D$ and  the direction of $f'$. We
have
 $$\mathscr D_{f'}f'=\sum_{\alpha=1}^n\mathscr D_{f'} f'_\alpha \partial_\alpha,$$
where
 $$\mathscr D_{f'} f'_\alpha=\sum_\beta f'_\beta \mathscr D_\beta f'_\alpha= \partial_zf'_\alpha+
 \sum_{\beta,\gamma}{\Gamma}_{\beta\gamma}^\alpha
 f'_\beta f'_\gamma.$$
Then, the holomorphic mapping
 $$f'\wedge\cdots\wedge
 f^{(n)}: \ \mathbb{C} \rightarrow E^{\frac{n(n-1)}{2}}\otimes K^*_X$$
is well defined.  In particular, we
obtain a holomorphic curve
$$
 f^{(2)}\wedge\cdots\wedge
 f^{(n)}: \ \mathbb{C} \rightarrow E^{\frac{n(n-1)}{2}}\otimes \bigwedge^{n-1}T_X.
$$The image of $f$ is called {\it autoparallel with respect to the
connection} $\mathscr D$ if
$$
  f'\wedge\cdots\wedge f^{(n)}\equiv 0.
$$

By means of meromorphic connections,  Siu \cite{Siu87}\index{Siu}  obtained  a defect relation.  A special case
of {\it Siu's Theorem}\index{Siu's Theorem} is presented by Nadel
\cite{Nadel}\index{Nadel} as follows
\begin{theorem}[Siu]\label{Siu-thm-Nst}
Let $f:\mathbb{C}\rightarrow X$ be a transcendental
holomorphic curve. If  $E^{-n(n-1)/2}\otimes K_X$ is ample, 
then  either $f(\mathbb{C})$ is  contained in the pole divisor $(t)$ 
of   $\mathscr D$ or $f(\mathbb{C})$ is 
autoparallel with respect to $\mathscr D$.
\end{theorem}
Nadel \cite{Nadel} remarked that the restriction ``\emph{transcedental}"  in Theorem \ref{Siu-thm-Nst} is not needed  since one may always replace $f(z)$ by $f(e^z).$ 
In what follows, we  extend Theorem \ref{Siu-thm-Nst}.  We first establish the following  Second Main Theorem type inequality: 
\begin{theorem}\label{mainthm2} 
Let $L$ be an ample line bundle over $X.$
Let $f:\mathbb{C}\rightarrow X$ be a 
holomorphic curve. 
Assume that  $f(\mathbb C)$ is not autoparallel with respect to  $\mathscr D.$
If $f(\mathbb C)\nsubseteq F^{-1}_\varphi(0)$ with  $\varphi= (f^{(2)}\wedge\cdots\wedge f^{(n)})_f,$  then
\begin{equation*}
T_f(r,K_X)+ N_{f, {\rm Ram}}(r)\leq_{\rm exc} \frac{n(n-1)}{2}T_f(r,E)+ O\Big(\log T_f(r,L)+\log r\Big).
\end{equation*}
\end{theorem}
\begin{proof} Since  $f(\mathbb C)$ is not autoparallel with respect to  $\mathscr D,$ then 
  $$ f'_f\wedge \big(f^{(2)}\wedge\cdots\wedge f^{(n)}\big)_f\not\equiv 0.$$
Equip $L$ with a Hermitian metric $h$ such that 
$\omega:=c_1(L,h)>0.$ Set $F_\varphi[\omega^n]=g^2f^*\omega.$ From  (6.3.17) in \cite{HY} (see also \cite{Hu1}), we note that 
 $$ \int_{|z|=r}\log g\frac{d\theta}{2\pi}\leq
 \int_{|z|=r}\log \|\varphi\|\frac{d\theta}{2\pi}\leq_{\rm exc} O\Big(\log T_f(r,L)+\log r\Big) .$$
So,  the proof   can be completed similarly to Theorem~\ref{mainthm}. 
 \end{proof}
 By Nadel's remark in above, Theorem~\ref{mainthm2}  yields immediately that 
 \begin{cor}\label{Siu-thm-Nst2}
Let $f:\mathbb{C}\rightarrow X$ be a  
holomorphic curve.
 If $E^{-n(n-1)/2}\otimes K_X$ is pseudo ample, 
then either $f(\mathbb{C})$ is  contained in the pole divisor $(t)$
of  $\mathscr D$ or $f(\mathbb{C})$ is 
autoparallel with respect to $\mathscr D$.
\end{cor}

In further, by the use of  techniques of  Siu \cite{Siu87} and Nadel \cite{Nadel},  the second possible statement  for the conclusion of Corollary \ref{Siu-thm-Nst2} can be also removed.
Namely, we have
 \begin{theorem}\label{Siu-thm-Nst21}
Let $f:\mathbb{C}\rightarrow X$ be a  
holomorphic curve.
 If $E^{-n(n-1)/2}\otimes K_X$ is pseudo ample, 
then  $f(\mathbb{C})$ is  contained in the pole divisor $(t)$
of  $\mathscr D.$ 
\end{theorem}

\vskip\baselineskip

\label{lastpage-01}
\end{document}